# Quasiconservation Laws and Suppressed Transport in Weakly Interacting Localized Models

### v1

@@ IV.3.1 Transport in the Single-Particle Anderson Model
Figure 9: Analysis of $\|\hat{M}^{(0)}\|_{F}$ for the noninteracting Anderson Model for 10000 disorder realizations, open boundary conditions, and with $i_{0}=\lfloor L/2\rfloor$. (a)–(d): Calculated probability density functions of $\|\hat{M}^{(0)}\|_{F}$. Each panel represents different disorder strengths $W=0.6,1.6,2.6,3.1$. Within each panel, each colored curve represents different system sizes from $L=20$ to $L=100$ in steps of $10$. The distributions are converged at large $L$, consistent with the presence of localization for arbitrarily weak disorder. (e): Medians of the distributions in (a)-(d), $\mathrm{med}\|\hat{M}^{(0)}\|_{F}$, as a function of $L$. Each curve represents a different disorder strength. Here, one sees that the curves saturate at large $L$, reflecting that, in the thermodynamic limit, charge cannot freely flow in the single-particle Anderson model due to localization. In the no disorder case $W=0$, the median scales as $\approx\frac{1}{4}\sqrt{L}$.
@@ IV.3.2 Transport in the Single-Particle Quasiperiodic Model
Figure 11: Analysis of the Frobenius norm of the charge transport capacity, $\|\hat{M}^{(0)}\|_{F}$, for the noninteracting quasiperiodic model with open boundary conditions. The analysis is done with 10000 evenly spaced phase shifts $\delta$ between $\delta=0$ and $\delta=\frac{1}{k}$, with $i_{0}=\lfloor L/2\rfloor$, and wavenumber $k=\sqrt{2}$. (a)–(d): Calculated probability density functions of $\|\hat{M}^{(0)}\|_{F}$. Within each panel, each colored curve represents a different system size, going from $L=20$ to $L=100$, in steps of $10$. Panels (a) and (b) are in the delocalized phase for $h=0.6$ and $1.6$, where one sees narrow and rightwards shifting peaks. This is a signature of unbounded transport in the delocalized phase. The localized phase is represented in panels (c) and (d) for potential depths $h=2.6,3.1$, which shows distributions converging sharply with increasing $L$. This reflects that transport is finite and limited in the localized phase. (e): We show the median of $\|\hat{M}^{(0)}\|_{F}$, $\mathrm{med}\|\hat{M}^{(0)}\|_{F}$, as a function of $L$. Each differently colored curve represents a different potential depth $h$, from $h=0.1$ to $h=3.1$ in steps of 0.5. We also show the curve at the transition $h=2$ (yellow). The curves in the extended phase $h<2$ show a square root scaling similar to that in the no-disorder case, while the curves in the localized phase for $h>2$ saturate to finite values at large $L$.
@@ IV.4.1 Anderson Model
Figure 12: Analysis of the Frobenius norm of the first-order correction to the charge transport capacity, $\|\hat{M}^{(1)}\|_{F}$, for the weakly-interacting Anderson Model across the link with $i_{0}=\lfloor L/2\rfloor$ and open boundary conditions. (a)–(d): Calculated probability density functions of $\ln(\|\hat{M}^{(1)}\|_{F})$ over 1000 disorder realizations. Each panel shows a different disorder strength $W=0.6,1.6,2.6,3.1$. Within each panel, each colored curve represents a different system size, with system sizes from $L=20$ to $L=100$ in steps of 10. The inset in (d) shows a distribution of $\|\hat{M}^{(1)}\|_{F}$ at $h=3.1$ with a natural log scale on the y and x-axes, with a $\|\hat{M}^{(1)}\|_{F}^{-2}$ (red line) comparison to the tail— see the text surrounding Eq. (21) for details on the form plotted. The convergence of the distributions at large disorder and $L$ indicates that the corrections to transport are bounded in the thermodynamic limit, suggesting that charge transport is suppressed in the presence of interactions at first-order in perturbation theory. (e): Plot of $\mathrm{med}\|\hat{M}^{(1)}\|_{F}$ as a function of $L$. Each curve represents a different disorder strength $W$, from $W=0.6$ to $3.1$ in steps of 0.5. The inset in (e) shows the same data but plotted using a log-log scale. At large $L$, the transport is bounded, indicating the presence of bottlenecks even when interactions are added.
Figure 13: Same data of $\mathrm{med}\|\hat{M}^{(1)}\|_{F}$ as in Fig. 12(e), but with the x-axis scaled roughly by the participation ratio ($\mathrm{PR}\sim W^{-1.78}$, see Appendix E for details), and y-axis scaled by a phenomenological scaling $W^{-5}$. The collapse of the curves here suggests that if one analyzes the corrections to large enough $L$, even the low disorder cases will have $\mathrm{med}{\|\hat{M}^{(1)}\|_{F}}$ saturate.
@@
To check this, we numerically calculate whether or not transport corrections larger than the median at a site $i_{0}$ will continue to stay larger than the median as one travels away from $i_{0}$. Thus, we define $\Theta_{i}\equiv\Theta\left(\|\hat{M}_{i}^{(1)}\|_{F}-\mathrm{med}\|\hat{M}_{i}^{(1)}\|_{F}\right)$. Here, $\Theta$ refers to the Heaviside function—thus, $\Theta_{i}$ is 1 if $\|\hat{M}_{i}^{(1)}\|_{F}$ is larger than the median of the $\|\hat{M}_{i}^{(1)}\|_{F}$ distribution, and 0 otherwise. We calculate the correlation coefficient $G(D)$ between $\Theta_{L/2}$ and $\Theta_{L/2+D}$ (for concision, we denote $\lfloor L/2\rfloor=L/2$, which is true in this case since we only perform calculations on lattices of even sizes), and limit ourselves to considering cases, out of $500$ disorder realizations, where $\|\hat{M}_{L/2}^{(1)}\|_{F}$ are large charge transport corrections (defined as corrections greater than the median of $\|\hat{M}_{L/2}^{(1)}\|_{F}$ for those 500 disorder realizations) at the central link $i_{0}=\lfloor L/2\rfloor$. The full expression for this correlation function is
@@ IV.4.2 Quasiperiodic Model
Figure 15: Numerical computations of the Frobenius norm of the first-order corrections to the charge transport capacity, $\|\hat{M}^{(1)}\|_{F}$, for the weakly-interacting quasiperiodic model with $i_{0}=\lfloor L/2\rfloor$ and open boundary conditions. Each calculation was done for over 500 phase shifts $\delta$ uniformly spaced between 00 and $1/k$. (a)–(d): Calculated probability density functions of $\ln(\|\hat{M}^{(1)}\|_{F})$. Each panel shows a different potential depth $h=0.6,1.6,2.6,3.1$. Within each panel, each colored curve represents a different system size, between $L=20$ and $L=100$, in steps of 10. The inset in (d) shows a distribution of $\|\hat{M}^{(1)}\|_{F}$ at $h=3.1$ with a log-scale on the y and x-axes, with a $\|\hat{M}^{(1)}\|_{F}^{-1.5}$ (red line) comparison to the tail— see the text surrounding Eq. (21) for details on the form plotted. (e): Plot of $\mathrm{med}\|\hat{M}^{(1)}\|_{F}$ as a function of $L$. Each curve represents a different potential depth $h$, from $h=0.6$ to $3.1$ in steps of 0.5. The inset in (e) shows the same plot but with a log-log scale.
@@ IV.5 On the Avalanche Instability at Weak Disorder
We note that a limitation of this work is that it cannot detect any avalanche instability. At low order in perturbation theory, an ergodic seed can only delocalize neighboring regions, preventing the thermal region from spreading throughout the system. From this perspective, avalanche instability is a fundamentally nonperturbative phenomenon. Recent work in Ref. [36] provided numerical evidence that the Anderson model, perturbed by weak interactions, immediately delocalizes up to a critical disorder $W^{*}$ that corresponds to our $W$ between 4 and 6. If avalanches indeed destabilize the localization at low $W$, our analysis did not detect such a transition. In the case of the Anderson model, our expressions for the LIOM corrections and charge transport capacity do not leave any room to capture a transition behavior akin to avalanches. The expressions for the AGP, LIOM corrections, and charge transport capacity operators, as in Eq. (17), Eq. (20), and Eq. (IV.2), only depend on the exponentially localized single-particle orbitals and energy eigenvalues. For the single-particle Anderson model, the energy spectrum and wavefunctions have similar properties at any disorder. Indeed, even at high disorder $W=6$, we verified that the quantities we calculated show the same qualitative behavior. In particular, the distributions still show the same slow power-law tails, and the distributions of the LIOM corrections and transport capacity converge at large enough $L$ for any disorder strength $W$. In conclusion, at low disorder our perturbative treatment may fail to detect the avalanche instability, while it is expected to be valid at large disorder.
@@ V Conclusion
Another open question our analysis can help address is the existence of MBL in $d>1$ dimensions. The seminal work of Ref. [61] argued that avalanches will inevitably destroy MBL in 2D and above. Despite this, experiments [16, 77] and numerics [78, 79, 80] have uncovered signatures of localization in 2D similar to those in the 1D case. Furthermore, the avalanche arguments reconcile poorly with the perturbative arguments presented in Ref. [31], which predicted the existence of MBL in any spatial dimension. It is fair to say that the existence of MBL in $d>1$ is still contested. However, MBL in 2D is extremely difficult to study on all fronts. Signatures obtained in experiments can be difficult to distinguish from glassy phenomena, which can have similar experimental signatures [81]. Furthermore, numerical studies have been largely impossible due to unfavorable scaling with system sizes. The analytical framework presented in this work can, in principle, be applied to any number of dimensions. In particular, one may ask if this approach can be used to study quasiconserved quantities in interacting models in 2D and above with nearest-neighbor interactions. The advantage of our method is the fact that larger system sizes are attainable (relative to what is achievable with exact diagonalization), and calculations scale as a polynomial with the number of sites $L$, as $O(L^{4})$. In this work, we calculated quantities up to 100 sites (which could correspond to a $10\times 10$ lattice in 2D), but it is possible to go to larger sizes if one has the computational resources and time. It would be extremely interesting to compare the behavior of quasiconserved quantities in the 2D case to those in the 1D case, as this may shed light on open questions about MBL in 2D. The generalization of the charge transport capacity operator $\hat{M}$ to higher dimensions is nontrivial, but one could study a similar object where the current is defined through a general $(d-1)$-dimensional surface cutting the $d$-dimensional lattice into two partitions. In 1D, this 0-dimensional surface is simply a single link, which was studied in detail in this work. In 2D, one would look at the current through a 1D line. Such a higher-dimensional charge transport capacity object could be an interesting probe for localization in higher dimensions. If one finds that the first-order corrections to LIOMs in 2D are not finite or the charge–transport–capacity–type arguments for bottlenecks do not work, that could strongly hint that localization indeed breaks down in 2D.
@@ Appendix C Additional Results for the AGP
Figure 16: Calculated probability density functions of the logarithm of the Frobenius norm of the AGP, $\ln\|\hat{X}\|_{F}$, similar to Fig. 1 (repeating panel with $W=2.1$ for easy comparisons here), for the weakly-interacting Anderson model. Each distribution was generated for 500 disorder realizations with open boundary conditions. Each panel represents a different disorder strength $W$, where disorders from $W=0.6$ to $W=3.1$ are shown, in steps of $0.5$. Within each panel, each curve represents a distribution for a different system size. Here, we show $L=20$ to $L=100$ in steps of $10$.
@@ D.1 Uniform Distribution of Eigenvalues
Figure 18: Probability density function of $x\equiv 1/(\epsilon_{\alpha}+\epsilon_{\beta}-\epsilon_{\gamma}-\epsilon_{\delta})$ calculated from Eq. (56). For this plot, $\epsilon_{\alpha}$ are assumed to have uniform distributions between $-1$ and $1$.
@@ D.2 Density of States of the Quasiperiodic Model
Figure 19: Calculated density of states of the quasiperiodic model, where onsite potential values are given by Eq. (4). Here, we fix $k=\sqrt{2}$ and $L=100$. The DOS is calculated over the whole spectrum of each of $\hat{H}$ with $500$ different phases $\delta$ between $\delta\in\left[0,\frac{1}{k}\right]$.
@@ Appendix E Participation Ratio Data
For an orbital with a large localization length in 1D, $\text{PR}\sim\xi$ 888 The participation ratio, while a proxy for the localization length, is not equivalent to it. The localization length $\xi$ is used to characterize the exponential decay of the orbitals $\phi_{i}$, where $\phi_{i}(j)\sim\exp[-|i-j|/\xi]$ On the other hand, the participation ratio measures the extent of the wavefunction $\phi$. Extracting $\xi$ directly from the PR is not straightforward—see, for instance, Ref. [84]. For our purposes, it is sufficient to look at the participation ratio rather than trying to directly calculate $\xi$..
For quick reference, in this section, we show data on how the median of the participation ratio grows with system size in Fig. 20(a), at different disorder strengths $W$. Here, the median is taken over $500$ disorder realizations, with the participation ratio being calculated for the eigenstate in the middle of the spectrum for each disorder realization. We are particularly interested in the behavior of the $\mathrm{PR}$ in the thermodynamic limit $L\rightarrow\infty$. At small $L$, the curves exhibit an initial power-law growth before saturating to a disorder-dependent plateau that decreases with increasing $W$. Moreover, the PR must vanish as $L\rightarrow 0$. These features motivate the use of the following fitting form for each $\mathrm{med[PR]}$

### v2

@@ IV.3.1 Transport in the Single-Particle Anderson Model
Figure 9: Analysis of $\|\hat{M}^{(0)}\|_{F}$ for the noninteracting Anderson Model based on 10000 disorder realizations, open boundary conditions, and with $i_{0}=\lfloor L/2\rfloor$. (a)–(d): Calculated probability density functions of $\|\hat{M}^{(0)}\|_{F}$. Each panel represents different disorder strengths $W=0.6,1.6,2.6,3.1$. Within each panel, each colored curve represents different system sizes from $L=20$ to $L=100$ in steps of 101010. The distributions are converged at large $L$, consistent with the presence of localization for arbitrarily weak disorder. (e): Medians of the distributions in (a)-(d), $\mathrm{med}\|\hat{M}^{(0)}\|_{F}$, as a function of $L$. Each curve represents a different disorder strength. Here, one sees that the curves saturate at large $L$, reflecting that, in the thermodynamic limit, charge cannot freely flow in the single-particle Anderson model due to localization. In the no disorder case $W=0$, the median scales as $\approx\frac{1}{4}\sqrt{L}$.
@@ IV.3.2 Transport in the Single-Particle Quasiperiodic Model
Figure 11: Analysis of the Frobenius norm of the charge transport capacity, $\|\hat{M}^{(0)}\|_{F}$, for the noninteracting quasiperiodic model with open boundary conditions. The analysis is done based on 10000 evenly spaced phase shifts $\delta$ between $\delta=0$ and $\delta=\frac{1}{k}$, with $i_{0}=\lfloor L/2\rfloor$, and wavenumber $k=\sqrt{2}$. (a)–(d): Calculated probability density functions of $\|\hat{M}^{(0)}\|_{F}$. Within each panel, each colored curve represents a different system size, going from $L=20$ to $L=100$, in steps of 101010. Panels (a) and (b) are in the delocalized phase for $h=0.6$ and 1.61.61.6, where one sees narrow and rightwards shifting peaks. This is a signature of unbounded transport in the delocalized phase. The localized phase is represented in panels (c) and (d) for potential depths $h=2.6,3.1$, which shows distributions converging sharply with increasing $L$. This reflects that transport is finite and limited in the localized phase. (e): We show the median of $\|\hat{M}^{(0)}\|_{F}$, $\mathrm{med}\|\hat{M}^{(0)}\|_{F}$, as a function of $L$. Each differently colored curve represents a different potential depth $h$, from $h=0.1$ to $h=3.1$ in steps of 0.5. We also show the curve at the transition $h=2$ (yellow). The curves in the extended phase $h<2$ show a square root scaling similar to that in the no-disorder case, while the curves in the localized phase for $h>2$ saturate to finite values at large $L$.
@@ IV.4.1 Anderson Model
Figure 12: Analysis of the Frobenius norm of the first-order correction to the charge transport capacity, $\|\hat{M}^{(1)}\|_{F}$, for the weakly-interacting Anderson Model across the link with $i_{0}=\lfloor L/2\rfloor$ and open boundary conditions. (a)–(d): Calculated probability density functions of $\ln(\|\hat{M}^{(1)}\|_{F})$ based on 1000 disorder realizations. Each panel shows a different disorder strength $W=0.6,1.6,2.6,3.1$. Within each panel, each colored curve represents a different system size, with system sizes from $L=20$ to $L=100$ in steps of 10. The inset in (d) shows a distribution of $\|\hat{M}^{(1)}\|_{F}$ at $h=3.1$ with a natural log scale on the y and x-axes, with a $\|\hat{M}^{(1)}\|_{F}^{-2}$ (red line) comparison to the tail— see the text surrounding Eq. (21) for details on the form plotted. The convergence of the distributions at large disorder and $L$ indicates that the corrections to transport are bounded in the thermodynamic limit, suggesting that charge transport is suppressed in the presence of interactions at first-order in perturbation theory. (e): Plot of $\mathrm{med}\|\hat{M}^{(1)}\|_{F}$ as a function of $L$. Each curve represents a different disorder strength $W$, from $W=0.6$ to 3.13.13.1 in steps of 0.5. The inset in (e) shows the same data but plotted using a log-log scale. At large $L$, the transport is bounded, indicating the presence of bottlenecks even when interactions are added.
Figure 13: Same data of $\mathrm{med}\|\hat{M}^{(1)}\|_{F}$ as in Fig. 12(e), but with the x-axis scaled roughly by the participation ratio ($\mathrm{PR}\sim W^{-1.78}$, see Appendix E for details), and the y-axis scaled by a phenomenological scaling $W^{-5}$. The collapse of the curves here suggests that if one analyzes the corrections to large enough $L$, even the low disorder cases will have $\mathrm{med}{\|\hat{M}^{(1)}\|_{F}}$ saturate.
@@
To check this, we numerically calculate whether or not transport corrections larger than the median at a site $i_{0}$ will continue to stay larger than the median as one travels away from $i_{0}$. Thus, we define $\Theta_{i}\equiv\Theta\left(\|\hat{M}_{i}^{(1)}\|_{F}-\mathrm{med}\|\hat{M}_{i}^{(1)}\|_{F}\right)$. Here, $\Theta$ refers to the Heaviside function—thus, $\Theta_{i}$ is 1 if $\|\hat{M}_{i}^{(1)}\|_{F}$ is larger than the median of the $\|\hat{M}_{i}^{(1)}\|_{F}$ distribution, and 0 otherwise. We calculate the correlation coefficient $G(D)$ between $\Theta_{L/2}$ and $\Theta_{L/2+D}$ (for concision, we denote $\lfloor L/2\rfloor=L/2$, which is true in this case since we only perform calculations on lattices of even sizes), and limit ourselves to considering cases, out of 500500500 disorder realizations, where $\|\hat{M}_{L/2}^{(1)}\|_{F}$ are large charge transport corrections (defined as corrections greater than the median of $\|\hat{M}_{L/2}^{(1)}\|_{F}$ for those 500 disorder realizations) at the central link $i_{0}=\lfloor L/2\rfloor$. The full expression for this correlation function is
@@ IV.4.2 Quasiperiodic Model
[…]
@@ IV.5 On the Avalanche Instability at Weak Disorder
We note that a limitation of this work is that it cannot detect any avalanche instability. At low order in perturbation theory, an ergodic seed can only delocalize neighboring regions, preventing the thermal region from spreading throughout the system. From this perspective, avalanche instability is a fundamentally nonperturbative phenomenon. Recent work in Ref. [36] provided numerical evidence that the Anderson model, perturbed by weak interactions, immediately delocalizes up to a critical disorder $W^{*}$ that corresponds to our $W$ between 4 and 6. If avalanches indeed destabilize the localization at low $W$, our analysis did not detect such a transition. In the case of the Anderson model, our expressions for the LIOM corrections and charge transport capacity do not leave any room to capture a transition behavior akin to avalanches. The expressions for the AGP, LIOM corrections, and charge transport capacity operators, as in Eq. (17), Eq. (20), and Eq. (IV.2), only depend on the exponentially localized single-particle orbitals and energy eigenvalues. For the single-particle Anderson model, the energy spectrum and wavefunctions have similar properties at any disorder. Indeed, even at high disorders $W=10$ and $W=20$, we verified that the quantities we calculated show the same qualitative behavior. In particular, the distributions still show the same slow power-law tails, and the distributions of the LIOM corrections and transport capacity converge at large enough $L$ for any disorder strength $W$. From the above calculation, it is clear that all the quantities calculated here would be saturated at very large $W$. Therefore, we focus on the small disorder regime for easier finite-size scaling analysis. In conclusion, at low disorder our perturbative treatment may fail to detect the avalanche instability, while it is expected to be valid at large disorder.
@@ V Conclusion
Another open question our analysis can help address is the existence of MBL in $d>1$ dimensions. The seminal work of Ref. [61] argued that avalanches will inevitably destroy MBL in 2D and above. Despite this, experiments [16, 81] and numerics [82, 83, 84] have uncovered signatures of localization in 2D similar to those in the 1D case. Furthermore, the avalanche arguments reconcile poorly with the perturbative arguments presented in Ref. [31], which predicted the existence of MBL in any spatial dimension. It is fair to say that the existence of MBL in $d>1$ is still contested. However, MBL in 2D is extremely difficult to study on all fronts. Signatures obtained in experiments can be difficult to distinguish from glassy phenomena, which can have similar experimental signatures [85]. Furthermore, numerical studies have been largely impossible due to unfavorable scaling with system sizes. The analytical framework presented in this work can, in principle, be applied to any number of dimensions. In particular, one may ask if this approach can be used to study quasiconserved quantities in interacting models in 2D and above with nearest-neighbor interactions. The advantage of our method is the fact that larger system sizes are attainable (relative to what is achievable with exact diagonalization), and calculations scale as a polynomial with the number of sites $L$, as $O(L^{4})$. In this work, we calculated quantities up to 100 sites (which could correspond to a $\times 10$ 10 lattice in 2D), but it is possible to go to larger sizes if one has the computational resources and time. It would be extremely interesting to compare the behavior of quasiconserved quantities in the 2D case to those in the 1D case, as this may shed light on open questions about MBL in 2D. The generalization of the charge transport capacity operator $\hat{M}$ to higher dimensions is nontrivial, but one could study a similar object where the current is defined through a general $(d-1)$-dimensional surface cutting the $d$-dimensional lattice into two partitions. In 1D, this 0-dimensional surface is simply a single link, which was studied in detail in this work. In 2D, one would look at the current through a 1D line. Such a higher-dimensional charge transport capacity object could be an interesting probe for localization in higher dimensions. If one finds that the first-order corrections to LIOMs in 2D are not finite or the charge–transport–capacity–type arguments for bottlenecks do not work, that could strongly hint that localization indeed breaks down in 2D.
@@ Appendix C Additional Results for the AGP
Figure 16: Calculated probability density functions of the logarithm of the Frobenius norm of the AGP, $\ln\|\hat{X}\|_{F}$, similar to Fig. 1 (repeating panel with $W=2.1$ for easy comparisons here), for the weakly-interacting Anderson model. Each distribution was generated absed on 500 disorder realizations with open boundary conditions. Each panel represents a different disorder strength $W$, where disorders from $W=0.6$ to $W=3.1$ are shown, in steps of 0.50.50.5. Within each panel, each curve represents a distribution for a different system size. Here, we show $L=20$ to $L=100$ in steps of 101010.
@@ D.1 Uniform Distribution of Eigenvalues
Figure 18: Probability density function of $x\equiv 1/(\epsilon_{\alpha}+\epsilon_{\beta}-\epsilon_{\gamma}-\epsilon_{\delta})$ calculated from Eq. (56). For this plot, $\epsilon_{\alpha}$ are assumed to have uniform distributions between 1-1- 1 and 111.
@@ D.2 Density of States of the Quasiperiodic Model
Figure 19: Calculated density of states of the quasiperiodic model, where onsite potential values are given by Eq. (4). Here, we fix $k=\sqrt{2}$ and $L=100$. The DOS is calculated over the whole spectrum of each of $\hat{H}$ with 500500500 different phases $\delta$ between $\delta\in\left[0,\frac{1}{k}\right]$.
@@ Appendix E Participation Ratio Data
For an orbital with a large localization length in 1D, $\text{PR}\sim\xi$ 999 The participation ratio, while a proxy for the localization length, is not equivalent to it. The localization length $\xi$ is used to characterize the exponential decay of the orbitals $\phi_{i}$, where $\phi_{i}(j)\sim\exp[-|i-j|/\xi]$ On the other hand, the participation ratio measures the extent of the wavefunction $\phi$. Extracting $\xi$ directly from the PR is not straightforward—see, for instance, Ref. [88]. For our purposes, it is sufficient to look at the participation ratio rather than trying to directly calculate $\xi$..
For quick reference, in this section, we show data on how the median of the participation ratio grows with system size in Fig. 20(a), at different disorder strengths $W$. Here, the median is taken over 500500500 disorder realizations, with the participation ratio being calculated for the eigenstate in the middle of the spectrum for each disorder realization. We are particularly interested in the behavior of the $\mathrm{PR}$ in the thermodynamic limit $L\rightarrow\infty$. At small $L$, the curves exhibit an initial power-law growth before saturating to a disorder-dependent plateau that decreases with increasing $W$. Moreover, the PR must vanish as $L\rightarrow 0$. These features motivate the use of the following fitting form for each $\mathrm{med[PR]}$